OpenStudy (anonymous):

whats 3x-2y=4can please give me the quardants from 0 to -4

OpenStudy (andriod09):

what exactly do you need?

OpenStudy (anonymous):

i need the x and y for that equation given while x is the numbers 0 ,1,-1,-2 ,-3,-4

OpenStudy (andriod09):

okay then.

OpenStudy (anonymous):

plaese hurrryyyyyyyy

OpenStudy (andriod09):

so, then 0,-1, -2, -3, and -4 are both terms X and Y?

OpenStudy (andriod09):

if so, then what you do is fill in the equation with the apropriate terms.

OpenStudy (anonymous):

i no that but when i did it didnt match up like if x=4 then y =4 thats what i got for one of them but it didnt look rite

OpenStudy (andriod09):

\[3(x)-2(y)=4\] it should be kinda oposites. what class is this? like algrebra or pre-ap?

OpenStudy (anonymous):

algrbra

OpenStudy (andriod09):

\[(x)=0, -1, -2, -3, and -4; and (y)=0, -1, -2, -3, and -4 \] correct?

OpenStudy (anonymous):

okay lets say that x is o,-1,-2,-3, -4 then what is y

OpenStudy (andriod09):

\[3(0)-2(0)\neq4\]\[3(-1)-2(-1)=-5\]\[3(-2)-2(-2)=-10\]\[3(-3)-2(-3)=-15\]\[3(-4)-2(-4)=-20\] post the question on ther so i can read it please

OpenStudy (andriod09):

that would be the answers if both (x) and (y) are the same.

OpenStudy (anonymous):

thanks

OpenStudy (andriod09):

is that what you need?

OpenStudy (anonymous):

about but i got the point

OpenStudy (andriod09):

okay. your welcome. BW i gave you a medel for your teamwork.

OpenStudy (anonymous):

awww thanks lol idk how to give medals

OpenStudy (andriod09):

click on the button next to my medel that has a 0

OpenStudy (anonymous):

i did i do it right

OpenStudy (andriod09):

yes you did. Thank you.

OpenStudy (anonymous):

np